Answer the following please.

1. Isoniazid is available in 30 mg unscored tablets. 0.06 g has been prescribed po. You should administer how many unscored tablets?

2. The physician has prescribed 25 mg of Dilaudid HP for administration IM. The Dilaudid HP is available in an ampule containing 20 mg in 1 cc. In preparing the drug for administration, what volume should you withdraw from the ampule?

3. 0.3 g of Elixophyllin has been prescribed po. 150 mg of Elixophyllin is available in capsules. How many capsules should be administered?

4. The physician has prescribed 50 mg of coumadin po. The coumadin is available in scored tablets labeled 100 mg. How many scored tablets should you administer?

5. 1.2 mg Lanoxin has been prescribed po. 0.6 mg of Lanoxin is available in scored tablets. How many scored tablets should be administered?

6. You have available an ampule of Nitrostat labeled 500 mg in 1 cc. 0.8 g of Nitrostat has been prescribed by the physician IV. What volume of drug will you prepare?

7. The physician has prescribed 0.3 g of Clinoril po. The Clinoril is available in scored tablets labeled 600 mg. How many scored tablets should you administer?

8. 10 mg of Ativan has been prescribed po. 5 mg of Ativan is available in unscored tablets. How many unscored tablets should be administered?

9. Vistaril is available in a vial labeled 40 mg in 1 mL. 0.1 g of Vistaril has been prescribed IM. How many mL should you withdraw from the vial in preparing to administer the drug?

10. You have available an ampule of Nitrostat labeled 250 mg in 1 cc. 0.4 g of Nitrostat has been prescribed by the physician IV. What volume of drug will you prepare?

11. The physician has prescribed 0.075 g of Clinoril po. The Clinoril is available in scored tablets labeled 150 mg. How many scored tablets should you administer?

12. 20 mg of Vistaril has been prescribed IM. How many mL should be prepared using a vial labeled 15 mg in 1 mL?

13. Lanoxicaps are available in 0.4 mg tablets. 0.3 mg has been prescribed po. You should administer how many tablets?

14. Lasix is available in a vial labeled 300 mg per 2 mL. 0.2 g has been prescribed by the physician po. You should prepare how many mL of the drug for administration?

15. 18 mg of liquid Dilantin has been prescribed po by the physician. What volume of Dilantin should you administer if the Dilantin is available in a strength labeled 40 mg in 3 cc?
Answer #1

1) 1 gram = 1000 mg

0.06 g = 0.06 * 1000 = 60 mg

So 2 tablets of 30 mg should be administered.

2)20 mg in 1 cc.

1 mg in 1/20 cc

25 mg in 1/20 * 25 = 1.25 cc

3) 0.3 g = 0.3 * 1000 = 300 mg

Number of capsules = 300/150 = 2

4)Number of tablets = 50/100 = 1/2